12 GeV Upgrade
At 21 weeks into the Six-Month Shutdown and with six weeks remaining, 12 GeV installation work is progressing as planned. Testing and commissioning of the first C-100 cryomodule is ongoing, and the second C-100 cryomodule has been installed and testing has begun. All 12 GeV activities during the shutdown period are integrated into the lab’s site-wide schedule, which is progressed weekly.
In Hall A, installation is going well, except for a problem with the polarized proton target. There was a splice between superconducting wire and normal wire. Not enough cooling was provided for this splice, which was buried in epoxy, and the splice overheated. The target group is proposing to re-make the splice(s) with longer sections of wire and more exposure to coolant. This repair job may take two-three weeks and could delay the start of the g2p/GEp experiment by a week or two.
In Hall C, installation of a ninth quartz bar started. This bar is identical to the eight quartz bars that function as the main detector for the Q-weak experiment, and it will be used to sensitively measure backgrounds in the detector hut. The Q-weak liquid hydrogen target was recently re-assembled after summer maintenance work. A test cooldown of the target was initiated in preparation for resumption of Q-weak running in November. Disassembly of the 300-metric-ton Hypernuclear Kaon Spectrometer was started. This spectrometer, which is being removed to make room for construction of the Super High Momentum Spectrometer, will be stored in the Experiment Staging Building.
The Central Helium Liquefier auxiliary transformer was replaced. The router power outage at the Machine Control Center has been completed. The low-conductivity water system has been installed to the 500 KeV dump, and the separator LCW system was returned to normal operation. The test plan for recovery of the radiofrequency systems is underway, and ongoing SRF commissioning is in progress. Pre-certification checks for the injector Personnel Safety System have begun. The beam switchyard beam dump has been filled and turned on, and the newly installed waveguides have been checked for leaks.
The FEL DC gun was rebuilt in the record time of one week, using parts from two gun assemblies, and the rebuilt gun assembly has now been mounted to the FEL injector. The gun electrodes were re-polished to 1-micron surface finish, cleaned for ultra-high vacuum and then mounted to the Pt-implanted original FEL alumina insulators. In parallel, the semi-load-lock system that allows a cathode changeout without breaking the gun vacuum environment was assembled and is being tested this week. Larger sulfur hexafluoride enclosures needed to accommodate this system were brought from the FEL gun test stand to the FEL vault and adjusted to fit in with the existing SF6 FEL system.
JLab Calendar of Events
Oct. 7: United Way Kickoff Picnic at Quark Cafe
Environment, Safety, Health and Quality
Drive Safely Work Week: Oct. 3-7
Computing and Networking Infrastructure
Major Systems Outage on Friday, Dec. 23
United Way Campaign at JLab Kicks Off With Oct. 7 Picnic
A United Way fund-raising raffle for each of two 16 gigabyte iPad2 devices will also begin during lunch on Oct. 7. Raffle tickets will be $1 each and tickets may be purchased at the table in the CEBAF Center lobby, noon-1 p.m. on workdays through Friday, Oct. 28. Only one iPad2 may be won per person. All money collected through the picnic and raffle will go to the United Way to support community services.
Open Skies for Foreign Air Travel Approved
Flu Vaccinations Available
Tracking Thomas at JLab
Congratulations this week go to Gary Croke, Greg Adams, Jason Willoughby, Percy Harrell, Andy Kowalski and Dick Owen, who were the first to correctly identify the Sept. 28 location. Honorable mentions go to Chris Williamson, Ron Bartek, Warren Funk, Mark Wiseman, Tina Johnson, Doug Higinbotham, David Anthony, Harry Fanning and Tom Powers. Check out the Tracking Thomas webpage for a better view of his last location and this week's new mystery photo.